Charles is now 5 times older than Elliot.
In 2 years will be three times older than Elliott.
How old is Charles now?

Found 2 solutions by Theo, Alan3354:
Answer by Theo(13342)

let c = age of charles now.
let e = age of elliot now.
today, you get c = 5 * e
two years from now, you get c + 2 = 3 * (e + 2)
in the second equation, simplify to get c + 2 = 3 * e + 6
subtract 2 from both sides of that equation to get c = 3 * e + 4
you have c = 5 * e and you have c = 3 * e + 4
this means that 5 * e = 3 * e + 4
subtract 3 * e from both sides of that equation to get 2 * e = 4
solve for e to get e = 2.
today, c = 5 * e results in c = 10
two years from now, c + 2 = 3 * (e + 2) results in 12 = 3 * 4
the solution is confirmed to be good.
the solution is that charles is 10 years old now and elliot is 2 years old now.
in two years charles will be 12 and elliot will be 4.

Answer by Alan3354(69443)

Charles is now 5 times older than Elliot.
In 2 years will be three times older than Elliott.
----------------
1 time older --> 2x as old
2 times older --> 3x as old
3 times older --> 4x as old
4 times older --> 5x as old
5 times older --> 6x as old
etc.
How old is Charles now?
=====================
C = Charles' age
E = Elliot's
----
C = 6E
C+2 = 3(E+2) = 3E+6 ---> C = 3E+4
-----
6E = 3E+4
3E = 4
E = 4/3
C = 8
-----------
Not an all integer solution because 5 times older is NOT the same as 5 times as old.
--- Is 1 time older the same age? If so, why say older?
=======================
Not your fault.
People who don't know what they're talking about should not write math problems, or anything else.